# The Beilinson–Bloch conjecture for some non-isotrivial varieties over global function fields

### v2

@@ 6.1. Motivic cohomology groups of $Y_{\eta}$
Here $\widetilde{W}$ is a choice of smooth proper scheme over a nonempty open $U\subseteq\mathbb{P}^{1}_{k}$ such that $\widetilde{W}\times_{U}K\cong W$, and the colimit is over all nonempty open $U^{\prime}\subseteq U$. As explained in [Jannsen2007, page 20], the arithmetic étale cohomology receives a regulator map
@@
For $a$, $b$, $c\in{\mathbb{Z}}$, the choice of $\widetilde{W}$ likewise equips $V=H^{b}(W_{\overline{K}},{\mathbb{Q}}_{\ell}(c))$ with an action of the étale fundamental group $\pi_{1}(U,\overline{\eta})$, where $\overline{\eta}$ is a geometric generic point of $U$. We define the arithmetic Galois cohomology of $V$ via
@@
We showed that $\operatorname{TB}$ holds for $Y$, and for all closed fibers of $f$. Now [Jannsen2007, Theorem 5.3] yields the below result.
@@
    For $i-2j\notin\{0,1\}$, we have $H^{i}_{\mathcal{M}}(Y_{\eta},{\mathbb{Q}}(j))=0$. For $i=2j+1$, there is an isomorphism
    $H^{2j+1}_{\operatorname{ar}}(Y_{\eta},{\mathbb{Q}}_{\ell}(j))\cong H^{1}_{\operatorname{ar}}(G_{\eta},H^{2j+1}(Y_{\overline{K}},{\mathbb{Q}}_{\ell}(j))).$
In particular, Jannsen’s conjecture [Jannsen1990, Conjecture 12.18] holds for $Y_{\eta}$.
We remark that Murre’s conjecture [Murre1993, 1.4], which [Jannsen2007, Theorem 5.3] also implies for $Y_{\eta}$, was previously known for cubic threefolds [dAMS98, Proof of Theorem 4.4].
@@ 7. The intermediate Jacobian
$\operatorname{CH}^{2}(Y_{\eta})_{0}=\operatorname{CH}^{2}(\mathfrak{h}^{3}(Y_{K}))=\operatorname{CH}^{1}(A)_{0}\cong A(K)\otimes_{\mathbb{Z}}{\mathbb{Q}}.$
It immediately follows that the Beilinson–Bloch conjecture for $Y_{\eta}$ (proved in theorem 6.4) is equivalent to the Birch and Swinnerton–Dyer conjecture for $A$. The latter is in turn equivalent via [KT2003, Theorem on page 541] to the strong Birch and Swinnerton–Dyer conjecture for $A$, which asserts that the Tate–Shafarevich group of $A$ is finite, that the $L$-function of $A$ vanishes at $s=1$ to order $\operatorname{rk}A(K)$, and that the leading Taylor coefficient of the $L$-function satisfies the expected formula.
@@
    $L(A,s)=\frac{\det(1-\operatorname{Frob}_{q}q^{-s}|H^{4}(Y_{\overline{{\mathbb{F}}}_{q}},{\mathbb{Q}}_{\ell}(1)))}{(1-q^{1-s})^{2}}.$
@@
    The base change of $A$ to $\overline{K}$ has endomorphism algebra isomorphic to $\mathbb{Q}$, and is not isogenous to an abelian variety defined over $\overline{{\mathbb{F}}}_{q}$.
@@
    The conductor of $A$ is the sum in $\operatorname{Div}(\mathbb{P}^{1}_{{\mathbb{F}}_{q}})$ of the points over which $f:Y\to\mathbb{P}^{1}_{{\mathbb{F}}_{q}}$ is not smooth. In particular, $A$ has everywhere semistable reduction.
@@
Claim (i) is immediate from proposition 5.10 and the Betti numbers of cubic hypersurfaces. The “big” geometric monodromy of the Tate module of $A$ (propositions 5.5 and 5.6), together with the Tate conjecture for endomorphisms of abelian varieties ([Tate1966], [Faltings1986], [Zarhin1974a], [Zarhin1974b], [MB1985]), yields (ii).
Finally, the conductor of $A$ is computed directly from proposition 5.9, as follows. Fix $t\in\mathbb{P}^{1}_{{\mathbb{F}}_{q}}$ with $Y_{t}$ singular, and $\overline{t}\to t$ a geometric point. Since the computation of the conductor exponent at $t$ is entirely local, we will elide any discussion of paths, and simply fix a vanishing cycle $\delta\in H^{3}(Y_{\overline{\eta}},{\mathbb{Q}}_{\ell}(1))$ for $\overline{t}$. Propositions 5.5 and 5.8 imply that $\delta$ is nonzero. The inertia $I_{\overline{t}}$ acts tamely on the rational $\ell$-adic Tate module $V_{\ell}(A)\cong H^{3}(Y_{\overline{\eta}},{\mathbb{Q}}_{\ell}(1))^{\vee}$ by proposition 5.9, and the inertia-invariant subspace is the orthogonal complement of $\delta$, so the conductor exponent of $V_{\ell}(A)$ at $t$ is
@@
The Birch and Swinnerton–Dyer conjecture in positive characteristic is known for isotrivial abelian varieties [Milne1968, Theorem 3] and for abelian varieties of analytic rank zero (for Jacobians over global function fields, combining [Ulmer2011, Formula (2.9.3)] with the arguments of [Tate1966b, Section 4] shows that analytic rank upper bounds algebraic rank; the extension of this inequality to general abelian varieties is well-known folklore). The above proposition thus indicates that the conjecture for $A$ is “hard”, in that it is not implied by either of these criteria. This may be unsurprising, since we used the highly nontrivial Tate conjecture for cubic fourfolds in order to prove it.
@@
We explicitly compute the $L$-function of $A$ when $X$ is the Fermat cubic fourfold. Let $\Delta$ be the base locus of $f$, and for simplicity assume that $\Delta$ is isomorphic to a blowup of $\mathbb{P}^{2}$ at six points (which can be guaranteed after passing to a finite extension of ${\mathbb{F}}_{q}$). Using the blowup formula for motives and proposition 7.1, the $L$-function of $A$ can now be calculated directly from the zeta function of $X$. The zeta functions of Fermat varieties were expressed by Weil in terms of Jacobi sums [KS79, Section 2.3]. In particular, when $p\equiv 2$ modulo 3, $H^{4}(X_{\overline{{\mathbb{F}}}_{q}},{\mathbb{Q}}_{\ell}(2))$ is generated by algebraic cycles (loc. cit., Theorem 2.10), so that
$\det(1-\operatorname{Frob}_{q}q^{-s}|H^{4}(X_{\overline{{\mathbb{F}}}_{q}},{\mathbb{Q}}_{\ell}(1)))$ $=(1-q^{1-s})^{23},$
$\det(1-\operatorname{Frob}_{q}q^{-s}|H^{4}(Y_{\overline{{\mathbb{F}}}_{q}},{\mathbb{Q}}_{\ell}(1)))$ $=(1-q^{1-s})^{30},$
@@
Remark 7.3.
Combining remark 3.5 with theorem 6.5, one may be able to prove analogues of the results of this section for abelian varieties which appear in the motives of certain cubic fivefolds over ${\mathbb{F}}_{q}(t)$.
Remark 7.4.
We conclude by proving a case of the Tate conjecture. Suppose that $\operatorname{char}(k)\neq 2$. Let $F$ be the Fano surface of $Y_{\eta}$, let $\mathcal{F}\to\mathbb{P}^{1}_{k}$ be a flat projective morphism with generic fiber $F$, and let $V\to\mathcal{F}$ be a projective resolution of singularities of $\mathcal{F}$ (which exists by [CP2008, Theorem 2.1]). The Tate conjecture holds for $F$ [Roulleau2013, Theorem 1], and the Birch and Swinnerton-Dyer conjecture holds for $\operatorname{Alb}(F)$, so the Tate conjecture holds for divisors on $V$ by [Geisser2021, Theorem 1.1]. The hard Lefschetz theorem [Deligne1980, Théorème 4.1.1] in turn implies the full Tate conjecture for $V$.

### v3

@@ 6.1. Motivic cohomology of $Y_{\eta}$
Here $\widetilde{W}$ is a choice of smooth proper scheme over a nonempty open $U\subseteq\mathbb{P}^{1}_{k}$ such that $\widetilde{W}\times_{U}K\cong W$, and the colimit is over all nonempty open $U^{\prime}\subseteq U$. As explained in [50, Section 13.6], the arithmetic étale cohomology receives a regulator map
@@
For $a$, $b$, $c\in{\mathbb{Z}}$, the choice of $\widetilde{W}$ likewise equips $V=H^{b}(W_{\overline{K}},{\mathbb{Q}}_{\ell}(c))$ with an action of the étale fundamental group $\pi_{1}(U,\overline{\eta})$. We define the arithmetic Galois cohomology of $V$ via
@@
Since we assumed $\operatorname{TB}(X)$, theorem 6.4 implies that $\operatorname{TB}$ holds for $Y$, and for all closed fibers of $f$. Now [50, Theorem 13.6.3] yields the below result.
@@
[…]
    (6.3) $H^{2j-1}_{\operatorname{ar}}(Y_{\eta},{\mathbb{Q}}_{\ell}(j))\cong H^{1}_{\operatorname{ar}}(G_{K},H^{2j-2}(Y_{\overline{\eta}},{\mathbb{Q}}_{\ell}(j))).$
  3. (iii)
    For all $j$, the cycle class map
    $\operatorname{CH}^{j}(Y_{\eta})_{\hom}\otimes_{\mathbb{Q}}{\mathbb{Q}}_{\ell}\xrightarrow{\sim}H^{2j}(Y_{\overline{\eta}},{\mathbb{Q}}_{\ell}(j))^{G_{K}}$
    and Abel–Jacobi map
    $\operatorname{CH}^{j}(Y_{\eta})_{0}\otimes_{\mathbb{Q}}{\mathbb{Q}}_{\ell}\xrightarrow{\sim}H^{1}_{\operatorname{cont}}(G_{K},H^{2j-1}(Y_{\overline{\eta}},{\mathbb{Q}}_{\ell}(j)))$
    are both isomorphisms.
In particular, Jannsen’s conjecture [47, Conjecture 12.18] holds for $Y_{\eta}$.
Note that claim (iii) of the theorem also follows from propositions 3.7 and 4.14. We remark that Murre’s conjecture [83, 1.4], which [50, Theorem 13.6.3] also implies for $Y_{\eta}$, was previously known for cubic threefolds [25, Proof of Theorem 4.4].
@@ 7. The intermediate Jacobian
$\operatorname{CH}^{2}(Y_{\eta})_{0}=\operatorname{CH}^{2}(\mathfrak{h}^{3}(Y_{\eta}))=\operatorname{CH}^{1}(A)_{0}\cong A(K)\otimes_{\mathbb{Z}}{\mathbb{Q}},$
where the last isomorphism is induced by a choice of polarization of $A$. It immediately follows that the Beilinson–Bloch conjecture for $Y_{\eta}$ (proved in theorem 6.4) is equivalent to the Birch and Swinnerton-Dyer conjecture for $A$. The latter is in turn equivalent via the main theorem of [57] to the strong Birch and Swinnerton-Dyer conjecture for $A$, which asserts that the Tate–Shafarevich group of $A$ is finite, that the $L$-function of $A$ vanishes at $s=1$ to order $\operatorname{rk}A(K)$, and that the leading Taylor coefficient of the $L$-function satisfies the expected formula.
@@
    $L(A,s)=\frac{\chi(H^{4}(Y_{\overline{k}},{\mathbb{Q}}_{\ell}(1)),s)}{(1-q^{1-s})^{2}}.$
@@
    The base change of $A$ to $\overline{K}$ has endomorphism algebra isomorphic to $\mathbb{Q}$, and is not isogenous to an abelian variety defined over $\overline{k}$.
@@
    The conductor of $A$ is the sum in $\operatorname{Div}(\mathbb{P}^{1}_{k})$ of the points over which $f:Y\to\mathbb{P}^{1}_{k}$ is not smooth. In particular, $A$ has everywhere semistable reduction.
@@
Claim (i) is immediate from proposition 5.11 and the Betti numbers of cubic hypersurfaces. The “big” geometric monodromy of the Tate module of $A$ (propositions 5.6 and 5.7), together with the Tate conjecture for endomorphisms of abelian varieties ([92], [34], [104], [103], [79]), yields (ii).
Finally, the conductor of $A$ is computed directly from proposition 5.10, as follows. Fix $t\in\mathbb{P}^{1}_{{\mathbb{F}}_{q}}$ with $Y_{t}$ singular, and $\overline{t}\to t$ a geometric point. Since the computation of the conductor exponent at $t$ is entirely local, we will elide any discussion of paths, and simply fix a vanishing cycle $\delta\in H^{3}(Y_{\overline{\eta}},{\mathbb{Q}}_{\ell}(1))$ for $\overline{t}$. Propositions 5.6 and 5.9 imply that $\delta$ is nonzero. The inertia $I_{\overline{t}}$ acts tamely on the rational $\ell$-adic Tate module $V_{\ell}(A)\cong H^{3}(Y_{\overline{\eta}},{\mathbb{Q}}_{\ell}(1))^{\vee}$, by proposition 5.10, and the inertia-invariant subspace is the orthogonal complement of $\delta$, so the conductor exponent of $V_{\ell}(A)$ at $t$ is
@@
The Birch and Swinnerton-Dyer conjecture in positive characteristic is known for isotrivial abelian varieties [75, Theorem 3] and for abelian varieties of analytic rank zero (for Jacobians over global function fields, combining [96, Formula (2.9.3)] with the arguments of [93, Section 4] shows that analytic rank upper bounds algebraic rank; the extension of this inequality to general abelian varieties is well-known folklore). The above proposition thus indicates that the conjecture for $A$ is “hard”, in that it is not implied by either of these criteria. This may be unsurprising, since we used the highly nontrivial Tate conjecture for cubic fourfolds in order to prove it.
@@
We explicitly compute the $L$-function of $A$ when $X$ is the Fermat cubic fourfold. Let $\Delta$ be the base locus of $f$, and for simplicity assume that $\Delta$ is isomorphic to a blowup of $\mathbb{P}^{2}$ at six points (which can be guaranteed after passing to a finite extension of ${\mathbb{F}}_{q}$). Using the blowup formula for motives and proposition 7.1, the $L$-function of $A$ can now be calculated directly from the zeta function of $X$. The zeta functions of Fermat varieties were expressed by Weil in terms of Jacobi sums [58, Section 2.3]. In particular, when $p\equiv 2$ modulo 3, $H^{4}(X_{\overline{{\mathbb{F}}}_{q}},{\mathbb{Q}}_{\ell}(2))$ is generated by algebraic cycles defined over ${\mathbb{F}}_{q}$ (loc. cit., Theorem 2.10), so that
$\chi(H^{4}(X_{\overline{{\mathbb{F}}}_{q}},{\mathbb{Q}}_{\ell}(1)),s)$ $=(1-q^{1-s})^{23},$
$\chi(H^{4}(Y_{\overline{{\mathbb{F}}}_{q}},{\mathbb{Q}}_{\ell}(1)),s)$ $=(1-q^{1-s})^{30},$
@@
The results presented thus far in this section have analogues in higher dimension. Let $V$ be a smooth cubic sixfold over $k$, and let $g:W\to\mathbb{P}^{1}_{k}$ be a Lefschetz pencil for $V$. Then $W_{\eta}$ is a smooth cubic fivefold over $K$. By corollary 3.21, there exists an abelian variety $B$ over $K$ such that $\mathfrak{h}^{5}(W_{\eta})\cong\mathfrak{h}^{1}(B)(-2)$. Proposition 3.7 thus implies that $\operatorname{BB}^{3}(W_{\eta})$ is equivalent to the Birch and Swinnerton-Dyer conjecture for $B$. In particular, if $V$ satisfies the hypotheses of theorem 6.5, then BSD holds for $B$.
The proof of the next proposition is identical to that of proposition 7.1, and omitted.
Proposition 7.3.
The following are true.
  1. (i)
    The $L$-function of $B$ is
    (7.1) $L(B,s)=\frac{\chi(H^{6}(W_{\overline{k}},{\mathbb{Q}}_{\ell}(2)),s)}{(1-q^{1-s})^{2}}.$
  2. (ii)
    The base change of $B$ to $\overline{K}$ has endomorphism algebra isomorphic to $\mathbb{Q}$, and is not isogenous to an abelian variety defined over $\overline{k}$.
  3. (iii)
    The conductor of $B$ is the sum in $\operatorname{Div}(\mathbb{P}^{1}_{k})$ of the points over which $g:W\to\mathbb{P}^{1}_{k}$ is not smooth. In particular, $B$ has everywhere semistable reduction.
Example 7.4.
[…]
$\dim_{{\mathbb{Q}}_{\ell}}H^{6}(V_{\overline{k}},{\mathbb{Q}}_{\ell}(2))+1=88\leq\operatorname{rk}B(K)+2\leq\dim_{{\mathbb{Q}}_{\ell}}H^{6}(W_{\overline{k}},{\mathbb{Q}}_{\ell}(2))=110,$
where the + 1+\ 1 on the left comes from the self-intersection of an ample divisor on $Q$.
We conclude by proving a case of the Tate conjecture. Let $F$ be the Fano surface of $Y_{\eta}$, let $\mathcal{F}\to\mathbb{P}^{1}_{k}$ be a flat projective morphism with generic fiber $F$, and let $\rho:T\to\mathcal{F}$ be a resolution of singularities of $\mathcal{F}$, such that $\rho$ is a projective morphism. Such a resolution exists by [23, Theorem 2.1].
Theorem 7.5.
The Tate conjecture holds for $T$.
Proof.
The Tate conjecture holds for $F$ (see theorem 3.24), and the Birch and Swinnerton-Dyer conjecture holds for $A=\operatorname{Alb}(F)$, so the Tate conjecture holds for divisors on $T$ by [39, Theorem 1.1]. The hard Lefschetz theorem [29, Théorème 4.1.1] in turn implies the full Tate conjecture for $T$. ∎